A motor has coils with a resistance of 26 Ω and operates from a voltage of 244 V. When the motor is operating at its maximum speed, the back emf is 146 V. 

(a) Find the current in the coils when the motor is first turned on. 

 (b) Find the current in the coils when the motor has reached maximum speed. 

 (c) If the current in the motor were 7.0 A at some instant, what is the back emf at that time? 
Answer #1

a)

V = 244 Volts

R = 26 ohm

i = V/R = 244/26 = 9.4 A

b)

at maximum speed :

Vb = 146

V = 244

hence i = (V - Vb)/R = (244 - 146)/26 = 3.77 A

c)

i = (V - Vb)/R

7 = (244 - Vb)/26

Vb = 62 volts